OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

subtract \(2x\) divide by \(3\)

OpenStudy (goldphenoix):

What is m - 2x? m - 2x = m -2x. There the same. Why? It's because you cannot combine them. (Not like term.) So your equation should look like: \[\large \large m -2x = 3y\] Now we want to find y. So we want to divide both side by 3. This would give us: \[\large \large \frac{ m-2x }{ 3 } =y\]

OpenStudy (goldphenoix):

Or: \[\large \large\frac{ 1 }{ 3 }m - \frac{ 2x }{ 3 } = y\]

OpenStudy (anonymous):

Solve M = 2x + 3y for y. is that for the literal equation? above

OpenStudy (goldphenoix):

What do you mean? the 2 equation I gave you is the answer. :) Just choose 1.